PARCO is considering a new project that complements its existing business. The machine required for the project costs $3.9 million. The marketing department predicts that sales related to the project will be $2.35 million per year for the next four years, after which the market will cease to exist. The machine will be depreciated down to zero over its four-year economic life using the straight-line method. Cost of goods sold and operating expenses related to the project are predicted to be 25 percent of sales. PARCO also needs to add net working capital of $150,000 immediately. The additional net working capital will be recovered in full at the end of the project’s life. The corporate tax rate is 30 percent. The required rate of return for PARCO is 13 percent. Should PARCO proceed with the project?


1
Expert's answer
2021-02-01T07:48:13-0500

Net cash flow: "2.35*0.75*0.7=1.23375" $ million

"PV=-(3.9+0.15)+\frac{1.23375}{1+0.13}+\frac{1.23375}{(1+0.13)^2}+\frac{1.23375}{(1+0.13)^3}+\frac{1.23375}{(1+0.13)^4}+\frac{0.15}{1.13^4}=-0.288248" $ million

No, since the project is unprofitable